# TA-Cleaner: A Fine-grained Text Alignment Backdoor Defense Strategy for Multimodal Contrastive Learning

### v2

@@ 4.3 Analysis
The pos-/neg- temperature factor Since we employ fine-grained alignment of positive and negative subtexts with images, it is essential to consider the relationship between the model’s focus on positive and negative samples and the final defense performance. This relationship can be modulated by adjusting the temperature factors $t_{p}$ and $t_{n}$ in Eq. 1 and 2 to achieve different levels of attention to positive and negative samples. Specifically, the smaller the value of the temperature factor, the higher the attention received. As shown in Figure 4(b), we conducted ablation experiments with five different sets of temperature factors and found that when $t_{p}$ is higher than $t_{n}$, the model de-emphasizes negative samples, leading to an inability to fine-tune the distribution of text features in the feature space, thus failing to actively distance itself from poisoned image features and resulting in poorer defense performance. Furthermore, if both factors are the same and relatively large, the fine-grained optimization weight of the model decreases, leading to the defense performance decrea. We found that a $t_{n}$ of 0.3 yields strong defense performance, and when $t_{p}=0.3$, the impact on BA is minimal. Therefore, we ultimately adopt $t_{p}=t_{n}=0.3$ as our default setting.
@@
The number of fine-grained texts in every epoch Furthermore, since we do not perform text augmentation on all samples, but rather randomly select a subset of samples to implement TA-Cleaner in each iteration, we explored the impact of sample quantity, as illustrated in Figure 4(c). It can be observed that for simpler attacks like BadNet, only 50 samples are sufficient to significantly reduce the ASR after the first iteration, achieving extremely fast and optimal defense performance. For more complex new attack techniques, such as BadCLIP, only 2000 to 3000 samples are needed. This represents a very small training cost compared to the scale of the fine-tuning dataset (500K).
5 Conclusions
In this paper, we focus on fine-tuning defense strategies against backdoor attacks targeting MCL. Building upon CleanCLIP, we propose an optimized approach utilizing fine-grained augmentation based on positive and negative subtexts. This method significantly mitigates attack risks while preserving the model’s normal feature expression capabilities, achieving efficient defense.
@@ 5 Conclusions
Limitations Since our proposed TA-Cleaner primarily addresses backdoor attacks in the image modality, the defense performance against text modality attacks remains unknown. In the future, we will further explore comprehensive and efficient defense methods that are effective across various modalities.
Appendix A The backdoor attack on CLIP
Our defense targets mainly backdoor attacks based on data poisoning. The process of such backdoor attacks against CLIP is shown in Figure 1. Against the pre-trained large model, the adversary fine-tunes the pre-trained large model with the dirty dataset adulterated with poisoned samples so that the model learns the feature alignment between the triggers and the target attack text. During the inference phase, the poisoned model behaves with normal output in the face of clean samples, but once the input image contains triggers, the model behaves with the malicious output specified by the adversary.
Figure 1: The data-poisoning backdoor attacks on CLIP.
Appendix B The loss functions of CLIP and CleanCLIP
CLIP Inspired by mapping images and texts into a shared feature embedding space $\mathbb{R}^{d}$, CLIP enables the model to directly understand the semantic relationship between them. The CLIP framework involves two encoders: an image encoder $f_{I}:I\rightarrow\mathbb{R}^{d}$ and a text encoder $f_{T}:T\rightarrow\mathbb{R}^{d}$, which transform the image and text data into representations of dimension $d$. The model is pre-trained through contrastive learning, leveraging vast amounts of internet image-text pairs $\{I_{i},T_{i}\}_{i=1}^{N}$ to learn the associations between images and texts. Mathematically, given an image embedding $z_{i}^{I}=f_{I}(I_{i})$ and a text embedding $z_{i}^{T}=f_{T}(T_{i})$ for a pair $(I_{i},T_{i})$, the model is trained using a multimodal contrastive loss $loss_{Clip}$ to align the text and image representations, which is shown in follows:
$loss_{Clip}=-\frac{1}{2N}\left(\sum_{j=1}^{N}\log\left(\frac{\exp\left(\langle z^{I}_{j},z^{T}_{j}\rangle/\tau\right)}{\sum_{k=1}^{N}\exp\left(\langle z^{I}_{j},z^{T}_{k}\rangle/\tau\right)}\right)+\sum_{k=1}^{N}\log\left(\frac{\exp\left(\langle z^{I}_{k},z^{T}_{k}\rangle/\tau\right)}{\sum_{j=1}^{N}\exp\left(\langle z^{I}_{k},z^{T}_{j}\rangle/\tau\right)}\right)\right),$ (1)
[…]
CleanCLIP CleanCLIP introduces a self-supervised loss based on multimodal data augmentation, which fine-tunes a clean dataset to reduce the impact of backdoor models. Their self-supervised loss can be formulated as follows:
$loss_{SS}=-\frac{1}{2N}\left(\sum_{j=1}^{N}\log\left(\frac{\exp\left(\langle z^{I}_{j},z^{\hat{I}}_{j}\rangle/\tau\right)}{\sum_{k=1}^{N}\exp\left(\langle z^{I}_{j},z^{\hat{I}}_{k}\rangle/\tau\right)}\right)+\sum_{j=1}^{N}\log\left(\frac{\exp\left(\langle z^{T}_{k},z^{\hat{T}}_{k}\rangle/\tau\right)}{\sum_{k=1}^{N}\exp\left(\langle z^{T}_{k},z^{\hat{T}}_{j}\rangle/\tau\right)}\right)\right),$ (2)
where the $z^{\hat{I}}_{i}$ and $z^{\hat{T}}_{i}$ represent the feature embeddings of augmented image $\hat{I_{i}}$ and text $\hat{T_{i}}$, i.e., $z^{\hat{I}}_{i}=f_{I}(\hat{I_{i}})$, $z^{\hat{T}}_{i}=f_{T}(\hat{T_{i}})$. And the CleanCLIP finetuning loss can be summarized as:
$loss_{CClip}=\gamma_{1}\cdot loss_{Clip}+\gamma_{2}\cdot loss_{SS}.$ (3)
Appendix C The linear probe test
We evaluated TA-C using linear-probe methods on a series of datasets introduced by Kornblith et al. [2019] to investigate whether it negatively impacts the model’s usability and transfer performance. For this evaluation, we tested models subjected to BadNet poisoning and defenses, with a learning rate of $31e-$ 3 during linear probe training. The corresponding test results are shown in Table 1. As observed, we achieved test results comparable to CleanCLIP, indicating that we significantly reduced the ASR without compromising the model’s performance and transferability. And the detailed information of these datasets are introduced as follows:
CIFAR10 and CIFAR100, introduced by Krizhevsky and Sutskever [2009], are small-scale color image datasets for image classification and object recognition tasks. CIFAR10 comprises 60,000 32x32 pixel color images in 10 classes, with 6,000 images per class. The dataset is split into 50,000 training images and 10,000 test images. CIFAR100 contains 100 classes grouped into 20 superclasses. Each class has 600 images, with 500 training images and 100 testing images. The superclasses organize the 100 classes in a semantic hierarchy.
ImageNet1K, often referred to as ILSVRC2012, is a subset of the ImageNet dataset Deng et al. [2009]. It consists of approximately 1.28 million training images and 50,000 validation images, covering 1,000 classes. The dataset is widely used for large-scale image recognition tasks.
DTD (Describable Textures Dataset) Cimpoi et al. [2014] contains images of textures grouped into 47 categories based on a list of adjective-noun texture descriptions. Each category has 120 images, totaling 5,640 images. The dataset is designed for research in texture analysis and recognition.
STL-10 dataset is designed for developing unsupervised feature learning, deep learning, and self-taught learning algorithms. It consists of 10 classes of unlabeled and labeled images, including aircraft, bird, car, cat, deer, dog, horse, monkey, ship, and truck. The images are 96x96 pixels in size.
SVHN (Street View House Numbers) is a real-world image dataset for digit recognition Netzer et al. [2011], derived from Google Street View images. It contains over 600,000 digit images coming from a variety of house numbers in Google Street View images. The dataset is partitioned into 73,257 training images, 26,032 test images, and 531,131 extra training images.
Food-101 dataset consists of 101 food categories with 1,000 images per category Bossard et al. [2014]. The images were collected from a large food image dataset available on the Internet. The dataset is designed for food recognition and related tasks.
Oxford-IIIT Pet Dataset Parkhi et al. [2012] is a large-scale dataset of pet images with fine-grained annotations. It consists of 37 pet categories with 200 images per category. The images exhibit large variations in scale, pose, and lighting. The dataset is designed for tasks such as fine-grained classification and segmentation.
RenderedSST2 is designed to evaluate optical character recognition (OCR) capabilities in a sentiment analysis context Socher et al. [2013]. It transforms textual sentiment labels into visual representations and comprises 60,000 32x32 rendered images of sentiment-labeled text (50,000 training images and 10,000 test images). The sentences belong to one of the 10 sentiment classes.
Num Train 50000 50000 50000 3760 5000 73257 75750 3680 6920
Num Test 10000 10000 50000 1880 8000 23062 25250 3669 1821
Num Classes 10 100 1000 47 10 10 101 37 2
Pre-trained 85.01 60.06 72.85 66.70 96.57 52.59 84.51 80.27 70.68
NoDefense 83.49 58.08 72.35 65.64 96.31 48.19 84.19 78.60 70.46
CleanCLIP 83.07 60.33 72.46 65.74 95.59 50.54 82.04 76.70 70.35
TA-C 83.32 60.01 72.37 65.78 95.83 51.06 83.01 77.21 70.35

### v3

@@ 4.3 Analysis
The pos-/neg- temperature factor Since we employ fine-grained alignment of positive and negative subtexts with images, it is essential to consider the relationship between the model’s focus on positive and negative samples and the final defense performance. This relationship can be modulated by adjusting the temperature factors $t_{p}$ and $t_{n}$ in Eq. 1 and 2 to achieve different levels of attention to positive and negative samples. The smaller the temperature factor, the higher the attention received. As shown in Figure 4(b), we conducted ablation experiments with five different sets of temperature factors and found that when $t_{p}$ is higher than $t_{n}$, the model de-emphasizes negative samples, leading to an inability to fine-tune the distribution of text features in the feature space, thus failing to actively distance itself from poisoned image features and resulting in poorer defense performance. Furthermore, if both factors are the same and relatively large, the fine-grained optimization weight of the model decreases, leading to the defense performance decrea. We found that a $t_{n}$ of 0.3 yields strong defense performance, and when $t_{p}=0.3$, the impact on BA is minimal. Therefore, we ultimately adopt $t_{p}=t_{n}=0.3$ as our default setting.
@@
The number of samples in every epoch applied CleanerCLIP Furthermore, since we do not perform text augmentation on all samples, but rather randomly select a subset of samples to implement CleanerCLIP in each iteration, we explored the impact of sample quantity, as illustrated in Figure 4(c). It can be observed that for simpler attacks like BadNet, only 50 samples are sufficient to significantly reduce the ASR after the first iteration, achieving extremely fast and optimal defense performance. For more complex new attack techniques, such as BadCLIP, only 2000 to 3000 samples are needed. This represents a very small training cost compared to the scale of the fine-tuning dataset (500K). Regarding model utility, we observe that increasing the number of neg-samples does not notably impact BA: when the num is 50, 250, 500, 750, 1000, 2000, and 3000, the Top-1 BA (%) are 51.53, 51.57, 51.51, 51.46, 51.47, 51.32, and 51.29, respectively. These results show that the increase in negative samples significantly reduces ASR while maintaining model utility, thanks to the stabilizing effect of positive sub-texts on clean sample features.
More ablation results In the supplementary materials, we provide additional experimental results to validate our effectiveness. We present the testing results on other MCL models in Sec 9, such as EVA-CLIP, and multimodal datasets like SBUCaption in Sec 10. Additionally, we provide a detailed comparison of the performance with CleanCLIP’s EDA text augmentation method in Sec 11, conducting self-supervised fine-tuning for each augmentation strategy to evaluate its defensive capabilities and compare them with our approach. In Sec 12, we also utilize existing open-source text augmentation strategies, like DeCLUTR, as a substitute for our counterfactual semantic enhancement component and compare their performance in fine-tuning defense.
5 Conclusion and limitation
In this paper, we focus on fine-tuning defense strategies against backdoor attacks targeting MCL. We propose CleanerCLIP, a counterfactual semantic enhancement method that effectively defends against backdoor attacks in multi-modal contrastive learning models. CleanerCLIP achieves superior defense performance across various datasets and attack scenarios, significantly reducing ASR while maintaining benign accuracy.
@@ 5 Conclusion and limitation
Limitations Since our proposed CleanerCLIP primarily addresses backdoor attacks in the image modality, the defense performance against text modality attacks remains unknown. In the future, we will further explore comprehensive and efficient defense methods that are effective across various modalities.